OpenStudy (anonymous):

You would like to have $170,000 in 12 years by making regular deposits at the end of each month in an annuity that pays 8.25% compounded monthly. How much of the $170,000 comes from interest? Round your answer to the nearest dollar.

jimthompson5910 (jim_thompson5910):

The first thing needed to be done is to find out the monthly payment

jimthompson5910 (jim_thompson5910):

so we use this formula P = FV( r/n )/((1+r/n)^(n*t) - 1 )

jimthompson5910 (jim_thompson5910):

P = monthly payment FV = future value r = annual interest rate n = compounding frequency t = # of years

jimthompson5910 (jim_thompson5910):

in this case P = unknown FV = 170000 r = 0.0825 n = 12 t = 12

jimthompson5910 (jim_thompson5910):

so plug all this in to get P = FV( r/n )/((1+r/n)^(n*t) - 1 ) P = 170000( 0.0825/12 )/((1+0.0825/12)^(12*12) - 1 ) P = 694.802351688371 P = 694.80

jimthompson5910 (jim_thompson5910):

so the monthly payment is $694.80 with me so far?

OpenStudy (anonymous):

yes. theres more to it?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you make deposits of $694.80 at the end of each month for 12*12 = 144 months

jimthompson5910 (jim_thompson5910):

so in total, you deposit 144*694.80 = 100,051.20 dollars over 12 years

jimthompson5910 (jim_thompson5910):

you deposit a total of 100,051.20 dollars but after 12 years, you will have $170,000 in the account so the difference is from the interest made

jimthompson5910 (jim_thompson5910):

170,000 - 100,051.20 = 69,948.80 So you made 69,948.80 dollars in interest

OpenStudy (anonymous):

Thank you so much! THis really helped

jimthompson5910 (jim_thompson5910):

glad it did

OpenStudy (anonymous):

I have another question!

jimthompson5910 (jim_thompson5910):

whats that

OpenStudy (anonymous):

You deposit $100 into a savings plan at the end of each year. The interest rate is 6% compounded annually. Find the value of the annuity after 7 years. Do not round until the final answer. Then, round to the nearest cent

jimthompson5910 (jim_thompson5910):

what you want is the future value (FV) of the annuity

jimthompson5910 (jim_thompson5910):

so you use this formula FV = P((1+r/n)^(n*t) - 1 )/( r/n ) in this case P = 100 r = 0.06 n = 1 t = 7

jimthompson5910 (jim_thompson5910):

I'll let you compute it. Tell me what you get.

OpenStudy (anonymous):

236.42?

OpenStudy (anonymous):

or is it 839.38?

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

FV = P((1+r/n)^(n*t) - 1 )/( r/n ) FV = 100((1+0.06/1)^(1*7) - 1 )/( 0.06/1 ) FV = 839.383764985601 FV = 839.38

jimthompson5910 (jim_thompson5910):

so that's how much will be in the account after 7 years

jimthompson5910 (jim_thompson5910):

Extra info: you made 7 payments of $100 each, so you put in 7*100 = 700 dollars total Interest Made: 839.38 - 700 = 139.38 dollars

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

np